# PANGAEA: A Global and Inclusive Benchmark for Geospatial Foundation Models

### v1

@@ A.10 AI4SmallFarms
The AI4SmallFarms dataset consists of 439,001 manually annotated agricultural field polygons distributed across 62 non-overlapping tiles, each approximately $5\times 5$ km, in Vietnam and Cambodia. The dataset includes multi-temporal 2entinel- image composites, created by aggregating monthly cloud-free images using median pixel values. The Sentinel-2 bands B2, B3, B4 and B8 are included for each tile. Additionally, high-resolution Google Map (GM) images are incorporated to improve boundary accuracy; however, these are not included in training the models, as the goal is to be able to perform farm boundary delineation from Sentinel-2 imagery. The digitization process involved detailed manual annotation from both Sentinel-2 and GM images, followed by a rigorous multi-step quality control procedure that included topological consistency checks and independent verification by multiple quality control teams.  Fig. A10 displays an example sample.
@@ A.10.1 Extensive Results
The results on the AI4SmallFarms dataset, presented in Table A17, show that the fully supervised UNet model significantly outperforms all evaluated GFMs. The UNet achieves an IoU of 46.34% and a mean F1 score of 62.36%, while the GFMs exhibit mIoU values ranging from 21.47% to 27.19% and m-f1 scores between 35.07% and 42.73%.

### v2

@@ A.10 AI4SmallFarms
The AI4SmallFarms dataset consists of 439,001 manually annotated agricultural field polygons distributed across 62 non-overlapping tiles, each approximately $5\times 5$ km, in Vietnam and Cambodia. The dataset includes multi-temporal Sentinel-2 image composites, created by aggregating monthly cloud-free images using median pixel values. The Sentinel-2 bands B2, B3, B4 and B8 are included for each tile. Additionally, high-resolution Google Map (GM) images are incorporated to improve boundary accuracy; however, these are not included in training the models, as the goal is to be able to perform farm boundary delineation from Sentinel-2 imagery. The digitization process involved detailed manual annotation from both Sentinel-2 and GM images, followed by a rigorous multi-step quality control procedure that included topological consistency checks and independent verification by multiple quality control teams.  Fig. A10 displays an example sample.
@@ A.10.1 Extensive Results
The results on the AI4SmallFarms dataset, presented in Table A17, show that the fully supervised UNet model significantly outperforms all evaluated GFMs. The UNet achieves an mIoU of 46.34% and a mean F1 score of 62.36%, while the GFMs exhibit mIoU values ranging from 21.47% to 27.19% and m-f1 scores between 35.07% and 42.73%.
@@ Appendix B Comparison with existing protocols
Appendix C Additional results
In this section of the Appendix, we present additional results that explore patterns between model performance and various characteristics of both pre-training and downstream datasets. Specifically, we investigate how performance varies across domains, geographic regions, and sensor resolutions, providing further insight into the factors that influence the generalization ability of geospatial foundation models.
Figure A12: Performance variation across domains for different GFMs. The heatmap illustrates relative performance deltas (higher is better) compared to the domain-specific average. Warmer colors indicate higher-than-average performance, while cooler colors denote underperformance. This visualization highlights the modality-task and domain-specific disparities across models, including performance gaps such as GFM-Swin’s relative drop in Flood and Marine tasks, potentially due to distributional shifts in pretraining.
Figure A13: Geographic variation in model performance across foundation models. This heatmap displays average performance scores across five geographic regions. Each cell reflects the model’s absolute performance within a specific region, highlighting spatial generalization capabilities. These patterns underscore the importance of geographic context in evaluating GFMs.
Figure A14: Performance variation of GFMs across different sensors and resolutions. This heatmap illustrates how model performance shifts depending on the source of input imagery, categorized by sensor types and their associated resolutions. The differences highlight the sensitivity of GFMs to input resolution and sensor characteristics.